# Nonequilibrium Phenomenology of Identified Particle Spectra in Heavy-Ion Collisions at LHC Energies

### v2

@@ Abstract
To determine the statistical operator $\rho(t)$ for a nonequilibrium process, we have to solve the Liouville-von Neumann equation $\frac{\partial}{\partial t}\rho(t)=\frac{i}{\hbar}[H,\rho(t)]$ with given initial conditions. These initial conditions are represented by the average values of a set of relevant observables $\{B_{i}\}$ in the past $t^{\prime}\leq t$ which characterize the state of the system. This information is used to construct the relevant statistical operator $\rho_{\rm rel}(t^{\prime})$ as the maximum of information entropy $S_{\rm inf}(t^{\prime})=-\langle\ln[\rho_{\rm rel}(t^{\prime})]\rangle_{\rm rel}$ under given constraints, the self-consistency conditions
@@
As well known from equilibrium statistics, these self-consistency conditions are taken into account via Lagrange parameters $\lambda_{i}(t^{\prime})$, and we obtain the generalized Gibbs distribution $\rho_{\rm rel}(t^{\prime})\propto\exp[-(\lambda_{i}(t^{\prime})B_{i})]$. The Lagrange parameters $\lambda_{i}(t^{\prime})$ must be eliminated using the self-consistency conditions (1) which are the nonequilibrium generalizations of the equations of state.
@@
A particular problem of the method of NSO is the selection of the set of relevant observables $\{B_{i}\}$. We have to consider long-living fluctuations because they need a long time to be built up dynamically and demand higher orders of perturbation theory. A minimum set of relevant observables consists of the conserved quantities: energy $H$ as well as the conserved numbers $N_{i}$. The Gibbs distribution $\rho_{\rm rel}(t)\propto\exp[-(H-\mu_{i}(t)N_{i})/T(t)]$ contains the Lagrange multipliers $\mu_{i}$, $T$ which are in general time dependent. They are not identical to the equilibrium parameters temperature and chemical potentials, but can be regarded as their generalizations in non-equilibrium.
@@
To describe the system formed in a heavy-ion collision, one must consider the fast system expansion. However, instead of introducing the collective velocity as another relevant observable and deriving the hydrodynamic equations from the NSO, we employ the blast-wave model parametrization in its standard form [7], assuming particles are emitted from the boost-invariant and cylindrically symmetric hypersurface at a constant value of the proper time $\tau=\sqrt{t^{2}-z^{2}}={\rm const}$. Then the resulting particle distribution function is given as:
@@
where $r$ is the transverse radius, $\phi$ the azimuthal angle in coordinate space, $\eta$ the space-time rapidity, $p_{T}$ the particle transverse momentum, $m_{T}$ the particle transverse mass, $\psi$ the azimuthal angle in momentum space, and $y$ is the particle rapidity. $T$ is the freeze-out temperature, $g_{i}$ and $\mu_{i}$ are the spin degeneracy factor and chemical potential of the $i$-th particle species with $a_{i}=+1$ for fermions and $a_{i}=-1$ for bosons. The velocity profile
is parameterized with a profile exponent $n$, transverse expansion velocity $\beta_{S}$, and the system radius $R$.
For resonances, we include additional mass attenuation in the form of the nonrelativistic Breit-Wigner distribution with constant width. Since particle-antiparticle ratios in the energy range of interest are approximately unity over a broad transverse momentum range, we set baryon, strangeness, and electric charge chemical potentials to zero. This reduces the number of free parameters while maintaining a reasonable fit quality. Then the particle chemical potential is defined as $\mu_{i}=\mu_{\pi}\delta_{i,\pi}+\mu_{K}\delta_{i,K}$.
@@
(anti-)protons [8]. For model B to be sensitive to changes in $\mu_{\pi}$, we extend the $p_{T}$ fit range for $\pi^{\pm}$ to $0.1-1.0$ GeV/c. To examine the overall fit quality, we introduce $\chi^{2*}$, which is the $\chi^{2}$ function evaluated for all spectra over the $p_{T}\in[0,2]$ GeV/c. It is remarkable that the fit with model B describes the data over a broad $p_{T}$ range, even outside of the fit ranges since $\chi^{2*}$ is $6.7$ times better in the extended model with finite effective chemical potentials. The NSO-based approximation B with non-equilibrium pion and kaon chemical potentials provides an excellent description of the particle spectra including the low-momentum enhancement.

### v3

@@ I Introduction
II The method of nonequilibrium statistical operator (NSO)
To determine the statistical operator $\rho(t)$ for a nonequilibrium process, we have to solve the Liouville-von Neumann equation
$\frac{\partial}{\partial t}\rho(t)=\frac{i}{\hbar}[H,\rho(t)]$ (1)
with given initial conditions. These initial conditions are represented by the average values of a set of relevant observables $\{B_{i}\}$ in the past $t^{\prime}\leq t$ which characterize the state of the system. This information is used to construct the relevant statistical operator $\rho_{\rm rel}(t^{\prime})$ as the maximum of information entropy
$S_{\rm inf}(t^{\prime})=-\langle\ln[\rho_{\rm rel}(t^{\prime})]\rangle_{\rm rel}$ (2)
under given constraints, the self-consistency conditions
@@ II The method of nonequilibrium statistical operator (NSO)
As well known from equilibrium statistics, these self-consistency conditions are taken into account via Lagrange parameters $\lambda_{i}(t^{\prime})$, and we obtain the generalized Gibbs distribution
$\rho_{\rm rel}(t^{\prime})=\dfrac{e^{-\sum_{i}\lambda_{i}(t^{\prime})B_{i}}}{{\rm Tr}e^{-\sum_{i}\lambda_{i}(t^{\prime})B_{i}}}.$ (4)
The Lagrange parameters $\lambda_{i}(t^{\prime})$ must be eliminated using the self-consistency conditions (3) which are the nonequilibrium generalizations of the equations of state.
@@
A particular problem of the method of NSO is the selection of the set of relevant observables $\{B_{i}\}$. We have to consider long-living fluctuations because they need a long time to be built up dynamically and demand higher orders of perturbation theory. A minimum set of relevant observables consists of the conserved quantities: energy $H$ as well as the conserved numbers $N_{i}$. The Gibbs distribution
$\rho_{\rm rel}(t)\propto\exp[-(H-\mu_{i}(t)N_{i})/T(t)]$ (6)
contains the Lagrange multipliers $\mu_{i}$, $T$ which are in general time dependent. They are not identical to the equilibrium parameters temperature and chemical potentials, but can be regarded as their generalizations in non-equilibrium.
@@
III Blast-wave model of particle freeze-out
To describe the system formed in a heavy-ion collision, one must consider the fast system expansion. However, instead of introducing the collective velocity as another relevant observable and deriving the hydrodynamic equations from the NSO [22], we employ the blast-wave model parametrization in its standard form [7], assuming particles are emitted from the boost-invariant and cylindrically symmetric hypersurface $\Sigma_{\mu}$ with the volume element $d\Sigma_{\mu}$ given by:
$d\Sigma_{\mu}=\left(\tau r\cosh\eta,0,0,-\tau r\sinh\eta\right)drd\phi d\eta,$ (7)
at a constant value of the proper time $\tau=\sqrt{t^{2}-z^{2}}={\rm const}$. Here $r$ is the transverse radius, $\phi$ is the azimuthal angle in coordinate space, and $\eta=\tanh^{-1}(z/t)$ is the space-time rapidity. In this model, $r\in\left[0;R\right]$ and $\eta\in\left[-\eta_{max};\eta_{max}\right]$, with $R$ and $\eta_{max}$ being the system size in transverse and longitudinal directions, respectively. The collective flow velocity in midrapidity is parametrized as
$u^{\mu}(r,\eta=0)=\left(\cosh\rho,\sinh\rho\cos\phi,\sinh\rho\sin\phi,0\right),$ (8)
where $\rho$ is the velocity profile:
with the profile exponent $n$ and transverse expansion velocity at the surface $\beta_{S}$. The full velocity field is obtained by boosting (8) in the $\eta$ direction. Having the freeze-out hypersurface and collective velocity fields defined, the particle invariant momentum distribution is given by the Cooper-Frye formula [23]:
[…]
where $g_{i}$ is the spin degeneracy factor of the $i$-th particle species, $p_{\mu}$ is the particle four-momentum, and $f_{i}(E)$ is the particle distribution function:
$f_{i}(E)=\left(\exp\left[\dfrac{E-\mu_{i}}{T}\right]+a_{i}\right)^{-1},$ (11)
with $a_{i}=+1$ for fermions and $a_{i}=-1$ for bosons, and $\mu_{i}$ and $T$ being the particle chemical potential and freeze-out temperature or their nonequilibrium generalizations from the NSO method if nonequilibrium is assumed. Then the resulting particle distribution function is given as:
@@ III Blast-wave model of particle freeze-out
where $p_{T}$ the transverse momentum of the particle, $m_{T}$ the transverse mass of the particle, $\psi$ is the azimuthal angle in momentum space and $y$ is the rapidity of the particle.
For resonances, we include additional mass attenuation in the form of the nonrelativistic Breit-Wigner distribution with constant width. Since particle-antiparticle ratios in the energy range of interest are approximately unity over a broad transverse momentum range, we set baryon, strangeness, and electric charge chemical potentials to zero. This reduces the number of free parameters while maintaining a reasonable fit quality. Then the particle chemical potential is defined as
$\mu_{i}=\mu_{\pi}\delta_{i,\pi}+\mu_{K}\delta_{i,K},$ (13)
where $\mu_{\pi}$ and $\mu_{K}$ are pion and kaon nonequilibrium effective chemical potentials.
@@ IV Results and discussion
(anti-)protons [8]. For model B to be sensitive to changes in $\mu_{\pi}$, we extend the $p_{T}$ fit range for $\pi^{\pm}$ to $0.1-1.0$ GeV/c. To examine the overall fit quality, we introduce $\chi^{2*}$, which is the $\chi^{2}$ function evaluated for all spectra over the $p_{T}\in[0,2]$ GeV/c. It is remarkable that the fit with model B describes the data over a broad $p_{T}$ range, even outside of the fit ranges since $\chi^{2*}$ is $6.7$ times better in the extended model with finite effective chemical potentials than in model A. The NSO-based approximation B with non-equilibrium pion and kaon chemical potentials provides an excellent description of the particle spectra including the low-momentum enhancement.
@@ V Conclusions
In a next step of the development of Zubarev’s NSO approach to describe the nonequilibrium evolution of the hadron resonance gas created in heavy-ion collisions would be the extension of the set of observables (3) to include the flow of particle species and to describe the formation of resonances as a result of interactions to be included to the model Hamiltonian. The goal of such an extended NSO approach would be to provide a consistent scenario for hadron production in heavy-ion collisions which shall justify the success of the thermal statistical model and its limitations. To give an example, the fugacity parameters that are introduced in thermal statistical models on an ad hoc basis to improve the description of experimental data (see, e.g. [15, 35]) could be related to the Lagrangian multipliers in the NSO formalism and would thus appear as the result of a consistent and systematic approach applied to describe the nonequilibrium system.